OpenStudy (shelbywyatt):

is 0 divided by 0 = n |n|≥o is this true?

OpenStudy (lncognlto):

0 divided by 0 is undefined doe...

OpenStudy (phi):

|n| >= 0 is true the | | sign means "make what is inside positive if it's negative" for example | -1 | = 1 if you already have a positive number, it does not do anything: | 1 | = 1 because | 0 | = 0 (0 is not negative, so the | | does not do anything) for any number you put inside the | |, you will get either 0 or a positive number we say that by writing | x | >= 0 or | n | >= 0

OpenStudy (skullpatrol):

$$\huge a \div a = \dfrac{a}{a} = a • \dfrac{1}{a} =1$$ Let a = 0 $$\huge 0 \div 0 = \dfrac{0}{0} = 0 • \dfrac{1}{0} = 0$$ because 0 times any number is 0, not 1. Therefore, division by 0 has no meaning in the set of real numbers.